OpenStudy (anonymous):

please help i will give medals find the axis of symmetry and the coordinates of the vertex of the graph of the function y = -3x^2 + 12x + 1

OpenStudy (anonymous):

@hoblos could you help??

OpenStudy (hoblos):

for a parabola with equation y = ax^2+bx + c the axis of symmetry is x = -b/2a can you get the axis now ?

OpenStudy (anonymous):

im not sure

OpenStudy (hoblos):

what are a,b and c in y = -3x^2 + 12x + 1 if you compare it to y = ax^2 + bx + c

OpenStudy (anonymous):

a = -3x^2 b = 12x c = 1

OpenStudy (anonymous):

im not sure how to put it together

OpenStudy (hoblos):

a,b and c are the coefficients i'll give you one , a= -3

OpenStudy (anonymous):

b= 12??

OpenStudy (anonymous):

like that

OpenStudy (hoblos):

correct and c ?

OpenStudy (anonymous):

c=1

OpenStudy (hoblos):

correctt so -b/2a = ?

OpenStudy (anonymous):

-3/2 = -1.5?

OpenStudy (hoblos):

where did you get the 3 and 2 from ?

OpenStudy (anonymous):

b = -3

OpenStudy (anonymous):

i don't know

OpenStudy (hoblos):

b=12 , and a=-3 so -b/2a = -12/2(-3) = ?

OpenStudy (hoblos):

\[-\frac{ 12 }{ 2 \times -3 }\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (hoblos):

you have to simplify it first

OpenStudy (anonymous):

-12/-6?

OpenStudy (hoblos):

correct then ?

OpenStudy (anonymous):

the answer is 2

OpenStudy (hoblos):

correct, so the axis of symmetry is x=2 now for the vertex the vertex is the intersection between the axis of symmetry and the parabola thus the abscissa of the vertex is x=2 replace x=2 in the equation what would be y ?

OpenStudy (anonymous):

y = 13??

OpenStudy (gahm8684):

i want a medal

OpenStudy (anonymous):

no